OpenStudy (anonymous):

Solve this quadratic equation. x2 + 5x + 3 = 0

OpenStudy (anonymous):

Tell me what is a, b and c here??

OpenStudy (anonymous):

1, 5 and 3

OpenStudy (anonymous):

So do you know how to find Discriminant??

OpenStudy (anonymous):

0_o NO

OpenStudy (anonymous):

Well, the formula is: \[D = b^2 - 4ac\] So you know a b and c just plug in to find D..

OpenStudy (anonymous):

D = 13

OpenStudy (callisto):

Since the question is solving the quadratic equation, you can apply the quadratic formula to solve x. \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Yes you are right.... Now to find roots: \[x = \frac{-b \pm \sqrt{D}}{2a}\]

OpenStudy (anonymous):